For a table manufacturing company, selling price for a table is $177.00 per Unit, Variable cost is $29.00 per Unit, rent is $3,391.00 per month and insurance is $288.00 per month. Company wants to expand its business and improve the table quality, it wants to increase the selling price for a table to $318.00 per Unit, Variable cost to $57.00 per Unit, bigger area will have rent $5,972.00 per month and insurance is $400.00 per month At what point will the company be indifferent between the current mode of operation and the new option?

A vendor prepares 100.00 hotdogs every day and sells at $18.00 /piece. For each hot dog, he spends $10.00 in the raw material. Additionally he spends $1.81 for packing each hotdog and monthly $52.00, $23.00, $8.00 as food truck rent, electricity and other expenses respectively. Lost sales are charged at $4.00 per lost sale. Leftover hotdogs can be sold for $3.On a particular day in June 105.00 people came wanting to buy a hotdog. Determine the vendor’s profit for that day? Assume there are 30 days in the month.

Please Please Please help!
Answer #1

For the table manufacturing company:

We are given the following information:

Particular Remark Original Expansion
SP Per unit $     171.00 $     253.00
VC Per unit $       29.00 $       47.00
Rent Per Month $ 4,171.00 $ 5,260.00
Insurance Per Month $     270.00 $     357.00

We need to find the monthly sales in units that will make the profits of each of these options equal and therefore the company will be indifferent between the two, so we need to equate the following equations to find the required number of units:

  • (SP - VC) x units = rent + insurance + profitoriginal
    • (171 - 29) x units = 4171 +270+ profitoriginal
  • (SP-VC) x units = rent + insurance + profiterpansion
    • 253 – 47) x units = 5260 +357 + profiterpansion
  • profiterpansion = profitoriginal
    • (171 – 29) x units – 4171 – 270 = (253 – 47) x units - 5260 - 357
    • units = 18.375
    • As furniture can't be in decimals, we have to round it up to next highest integer which is 19 units
  • To verify, we may calculate the profits under both the options with 19 units of goods sold
    • Particular Remark Original Expansion
      SP Per unit $      171.00 $      253.00
      VC Per unit $         29.00 $         47.00
      Rent Per Month $   4,171.00 $   5,260.00
      Insurance Per Month $      270.00 $      357.00
      Units sold Per Month 19.00 19.00
      Profit As per equation $ -1,743.00 $ -1,703.00
    • We can see that the difference in profits is not too high. if we use the exact number 18.375 then these two numbers are exactly the same
    • Particular Remark Original Expansion
      SP Per unit $      171.00 $      253.00
      VC Per unit $         29.00 $         47.00
      Rent Per Month $   4,171.00 $   5,260.00
      Insurance Per Month $      270.00 $      357.00
      Units sold Per Month 18.375 18.375
      Profit As per equation $ -1,831.75 $ -1,831.75
    • And the annual number is 18.375 x 12 = 220.50 units

Hotdog vendor:

We have calculated the profit in the below table:

Particular Remark Numbers Daily Number Explanation of Daily Number
Prepared hotdogs Given 100.000 100.00 Day sales of 100 hotdogs is given
SP Per unit $       18.00 $              18.00 remains constant
R/m Per unit $       11.00 $              11.00 remains constant
Packaging Per unit $         1.07 $                1.07 remains constant
Rent Per Month $       46.00 $                1.53 46/30 = 1.53
Electricity Per Month $       23.00 $                0.77 23/30 = 0.77
Other Exp Per Month 13.000 $                0.43 13/30 = 0.43
Customers visited Given 115.000 115.000 remains constant
Lost sales units 15.000 15.000 115-100 = 15
Charges of lost sales 90.000 90.000 150 x 6 = 90
Profit $            500.27 100 x (18-11-1.07)-1.53-0.77-0.43-90=500.27

So the daily profit is $500.27